# Multi-CATE: Multi-Accurate Conditional Average Treatment Effect Estimation Robust to Unknown Covariate Shifts

### v1

@@ 1 Introduction and Related Work
In these consequential domains, many firms/decision-makers face treatment decisions where other firms also need to make the same decision, although perhaps each with slightly different data distributions. For example, problems of clinical risk prediction, such as risk of a heart disease or medication treatment guidelines, are shared widely across hospitals, but each has its own distribution of patients in addition to idiosyncratic reporting, testing, and treatment patterns that can hinder external validity [Caruana et al., 2015]. Indeed, off-the-shelf, relatively simple clinical risk calculators developed on one population are often broadly deployed as a decision support tool in many locations, without the ability to share the originating individual-level data, or with data drift over time. In social settings, the Arnold Public Safety Assessment (PSA), trained on a proprietary dataset and used in hundreds of jurisdictions [Goel et al., 2021], is an example of a widely deployed tool. Its accompanying decision-making matrix is another example of a treatment recommendation rule made more widely available [Laura and Foundation, 2016], which can introduce disparities and poor treatment efficacy [Zhou, 2024]. Other examples include the design of algorithmic profiling in active labor market programs [Crépon and Van Den Berg, 2016, Bach et al., 2023, Körtner and Bonoli, 2023]: many different jurisdictions run different active labor market programs, and policymakers face questions about how to learn from what works elsewhere and how to scale up programs across heterogeneous locations.
A key challenge in these settings is to certify valid predictive performance of personalized causal effects for unknown deployment settings. For example, predictive risk calculators, such as those for chronic heart disease, learned on a specific population might induce biased estimation for different locales with different populations. As one example, the widely used Framingham risk score overestimates risk for Asian populations [Badawy et al., 2022]. This problem is not limited to earlier risk scores, but also modern ones: a sepsis predictive risk score provided by Epic, a major healthcare IT provider, fell short in a study of external validity on another population [Habib et al., 2021].
@@
On the other hand, we do want to leverage predictive information when it is available. How can we develop methods for heterogeneous treatment effect estimation so that a new hospital, without its own large database or in-house machine learning team, is still assured guarantees of low predictive bias on its own population, that might differ in unknown ways from a proprietary risk score that does not publish the original data?
@@
In this paper, we show how methods from multi-accurate learning [Hébert-Johnson et al., 2018, Kim et al., 2019] can endow conditional average treatment effect estimation with robustness to unknown covariate shifts. Indeed, the problem of confounding itself is a covariate shift problem, from the treated or control population to some reference population [Johansson et al., 2022]. Multi-accurate learning is a powerful and flexible framework that, by ensuring low predictive bias over a test function class, is also robust to combinations of these covariate shifts: those induced by confounding or unknown covariate shifts in the reference population. Although multi-calibrated and accurate learning originated from fairness motivations re: ensuring calibration/low prediction bias over rich subgroups, in this work we show how the adversarial test functions in the formulation also confer broad robustness against covariate shift. To highlight this flexibility, we use multi-accurate calibration on an extremely small clinical trial to correct a predictor from an confounded observational study.
@@
Though there is extensive work on establishing external validity and transportability of causal effects, most of this work assumes information about a target population. Drawing inspiration from Kim et al. [2022], which studied “universal adaptability” of estimating the ATE with bias robust to unknown covariate shifts, we learn CATE estimates that will maintain unbiased predictions under unknown target populations.
@@
Although causal inference and machine learning has witnessed significant methodological innovation either in orthogonal/statistical learning or other machine learning adaptations [Kennedy, 2023, Nie and Wager, 2020, Chernozhukov et al., 2018, Wager and Athey, 2018, Shalit et al., 2017, Hill, 2011], to name just a few, multi-accurate learning [Kim et al., 2019] introduces a different methodological toolkit related to boosting/adversarial formulations of conditional moment conditions. Recent advances in machine learning for conditional moment equations [Dikkala et al., 2020, Bennett and Kallus, 2023, Ghassami et al., 2022] typically develop min-max estimation algorithms that are unstable in practice; besides, the theory of conditional moment restrictions typically identifies finite-dimensional parameters rather than entire functions like the CATE. (See Section 5 for more extensive discussion of related work.) We conduct a thorough empirical study comparing finite- and large-sample performance of multi-accurate learning and other causal machine learning techniques more specifically tailored for causal structure. To summarize, we find that multi-accurate methods grant additional robustness against unknown covariate shifts while being competitive with more advanced causal machine learning methods in finite-samples. There is a robustness-efficiency tradeoff: the latter methods are designed to exploit in-distribution efficiency, which multi-accurate learning “off-the-shelf” does not. Nonetheless, our work connects these two previously unrelated lines of work and shows how multi-accurate learning “off-the-shelf” can address the problem of robust CATE estimation. Multi-accurate learning reduces prediction bias from model misspecification, just as is required for conditional average treatment effect estimation.
@@
In our thorough empirical study we find that our proposed multi-accurate T- and DR-learner perform well under unobserved covariate shift. Although our work does not suggest multi-accurate learning as a replacement for state-of-the-art causal machine-learning for in-distribution estimation, it does provide evidence that could inform further methodological improvements and variance reduction of multi-accurate learning for CATE estimation. In summary:
  • Multi-accurate learning can be used “off-the-shelf” to post-process CATE estimates based on differenced outcome regressions to endow them with robustness to unknown covariate shift.
  • Multi-accurate post-processing can improve CATE estimates with only black-box access to predictors and original data.
  • Alternative approaches to robustness against unknown shifts, like distributionally robust optimization, could change the robust-optimal predictor to a risk-sensitive one rather than the true CATE, but multi-accurate learning does not.
  • The multi-accuracy framework can approximate more advanced CATE estimators (such as the DR-learner [Kennedy, 2020, Semenova and Chernozhukov, 2021]) with appropriate selection of the test function class. That is, in Proposition 3 we show that multi-accurate post-processing of simple CATE estimates (T-learner) with a richer test function class can approximate a less-multi-accurate/less-robust but more-advanced CATE estimator, i.e. a multi-accurate DR-learner under a simpler test function class.
@@
The contributions of our work are the following. We propose multi-accurate post-processing of pseudo-outcome based CATE estimation to obtain unbiased prediction on unknown deployment populations. This approach can also flexibly adapt to a variety of covariate shifts from confounding to adversarial/unknown shifts: we illustrate by postprocessing a CATE estimator that combines large observational/small randomized data. We show in extensive experiments with simulations and real-world observational and randomized data from the Women’s Health Initiative how our approach achieves finite-sample gains in ensuring robust bias control (and correspondingly, MSE) under unknown distribution shifts.
@@ 2 Problem setup
We overview the problem setup and describe directly related prior work on multicalibration/multi-accuracy. See Section 5 for discussion of other methodological approaches.
Problem setup
@@
In different applications it will satisfy different assumptions, so we will later define different variants of $\mathcal{D}$. We first assume it arises from a randomized controlled trial or observational study under the assumption of weak ignorability, so that the following assumption about selection on unobservables holds.
@@
These are the typical so-called nuisance estimation functions used in common estimators. Sometimes we will refer to the true population functions as $\mu_{t}^{*},e_{t}^{*}$ to clarify.
@@
Later on, our work will focus on providing guarantees on conditional bias achieved by a CATE estimate $\hat{\tau}$ marginalized under a covariate distribution $X\sim Q_{X}$ that can be different from the distribution $X\sim P$ upon which the CATE estimate was trained:
@@
We write $Q_{X}(x),P_{X},P_{X_{1}},P_{X_{0}}$ for the marginal distribution of $X$ under $Q$, the marginal distribution of $X$ under $P$, and $P_{X_{1}},P_{X_{0}}$ under $X\mid T=1,X\mid T=0$ on the observed data, respectively. We also denote ${E}_{P}[\cdot],{E}_{P_{1}}[\cdot],{E}_{P_{0}}[\cdot]$ to denote marginalization over $X$ in the training data, $X\mid T=1,$ or $X=T=0$, respectively. For brief we write ${E}_{Q}[\cdot]$ to denote expectations under the unknown reference distribution $Q_{X}$ on $X$ (which can be extended to accommodate shifts beyond the typical covariate shift assumption in a slight abuse of notation). For example, $\mu_{t}\in\arg\min_{\mu}E_{P_{t}}[(Y-\mu)^{2}],$ e.g. by default, regression in each treated arm minimizes the MSE under the covariate distribution of each treatment arm.
Figure 1: Schematic of setting 1 (external shift), and setting 2 (learning from large observational and small RCT data)
@@ 2.1 Robustness to unknown deployment shifts
[…]
@@
We will call this type of unknown deployment shift an “external shift”. This is analogous to the unknown shift setting studied in Kim et al. [2022], as well as other literature on unknown covariate shifts [Jeong and Namkoong, 2020, Subbaswamy et al., 2021, Hatt et al., 2021]. In contrast to an extensive literature on transportability and external validity, we focus on the case of a-priori unknown deployment shifts.
If suitably nonparametric CATE estimation indeed recovered the Bayes-optimal predictor in finite samples, there would be no issue of unknown deployment shifts. But because in finite samples it generally does not, modifying estimation to protect against unknown deployment shifts can protect against misspecification and finite-sample issues. For example, misspecified CATE estimation is vulnerable to unknown covariate shift. The conventional mean-squared error MSE can be nonzero for the Bayes-optimal predictor $\mu_{1}^{*}(X)=\operatorname{E}[Y(1)\mid X]$. If the conditional bias or variance in $Y$ is heteroskedastic (i.e. varies in $x$), the prediction MSE changes as external shifts change the marginalizing covariate distribution. Later on we will use multi-accurate learning to post-process CATE estimates to ensure robustness against covariate shifts represented by a function class of likelihood ratios.
@@ 2.2 Unobserved confounding: observational data with RCT
We consider a different setting where unknown covariate shifts may arise: a large observational dataset and small randomized trial. The observational study may be subject to unobserved confounding. On the other hand, the sample size of the randomized data may be small, so that learning conditional causal effects solely from randomized data is unsupported. This regime is common in clinical settings, such as the parallel Women’s Health Initiative observational study and clinical trial [Machens and Schmidt‐Gollwitzer, 2003]; see also [Colnet et al., 2020, Yang et al., 2020] and [Bareinboim and Pearl, 2013] for identification results for the related setting of data fusion.
@@
Assume Assumption 2. Suppose an observational dataset $\mathcal{D}_{o}$, collected under violations of Assumption 1 (the observational data were collected under unobserved confounders), and a randomized dataset $\mathcal{D}_{r}$, where Assumption 1 holds (the randomized data are unconfounded).
@@
$|\mathcal{D}_{obs}|\gg|\mathcal{D}_{r}|$
Assumption 3 is not necessary for identification, but it describes the relevant regime where the method is helpful: if instead $|\mathcal{D}_{r}|\gg|\mathcal{D}_{o}|$, unbiased CATE estimation is possible from the randomized data alone.
3 Method
3.1 Background on Estimation
Conditional average treatment effect estimation
We briefly discuss a few options for estimating $\tau$, upon which we will build multicalibrated approaches in differing shift scenarios. The S-learner models outcome given covariates and treatment indicator $(x,t)$, that is, the covariate vector is simply augmented with a column for the treatment indicator. The corresponding CATE estimator imputes the counterfactual outcome:
$\hat{\tau}(x)=\hat{\mu}(x,1)-\hat{\mu}(x,0).$
The T-learner differences two regressions for the conditional means of Y for treated and untreated:

### v2

# Multi-Accurate CATE is Robust to Unknown Covariate Shifts

@@ 1 Introduction
In these consequential domains, many firms/decision-makers face treatment decisions where other firms also need to make the same decision, although perhaps each with slightly different data distributions. For example, problems of clinical risk prediction, such as risk of a heart disease or medication treatment guidelines, are shared widely across hospitals, but each has its own distribution of patients in addition to idiosyncratic reporting, testing, and treatment patterns that can hinder external validity [Caruana et al., 2015]. Indeed, off-the-shelf, relatively simple clinical risk calculators developed on one population are often broadly deployed as a decision support tool in many locations, without the ability to share the originating individual-level data, or with data drift over time. In social settings, the Arnold Public Safety Assessment (PSA), trained on a proprietary dataset and used in hundreds of jurisdictions [Goel et al., 2021], is an example of a widely deployed tool. Its accompanying decision-making matrix is another example of a treatment recommendation rule made more widely available [Laura and John Arnold Foundation, 2016], which can introduce disparities and poor treatment efficacy [Zhou, 2024]. Other examples include the design of algorithmic profiling in active labor market programs [Crépon and Van Den Berg, 2016, Bach et al., 2023, Körtner and Bonoli, 2023]: many different jurisdictions run different active labor market programs, and policymakers face questions about how to learn from what works elsewhere and how to scale up programs across heterogeneous locations.
A key challenge in these settings is to certify valid predictive performance of personalized causal effects for unknown deployment settings, each of which could have a different target covariate distribution. For example, predictive risk calculators, such as those for chronic heart disease, learned on a specific population might induce biased estimation for different locales with different populations. As one example, the widely used Framingham risk score overestimates risk for Asian populations [Badawy et al., 2022]. This problem is not limited to earlier risk scores, but also affects modern ones: a sepsis predictive risk score provided by Epic, a major healthcare IT provider, fell short in a study of external validity on another population [Habib et al., 2021].
@@
On the other hand, we do want to leverage predictive information when it is available. How can we develop methods for heterogeneous treatment effect estimation so that a new hospital, without its own large database or in-house machine learning team, is still assured guarantees of low predictive bias on its own population, that might differ in unknown ways from a proprietary risk score that does not publish the original data? Importantly, while some methods guarantee validity of prediction models for a known target distribution, our previous examples highlight important cases when we want to guarantee low prediction bias under many unknown target distributions.
@@
In this paper, we show how methods from multi-accurate learning [Hébert-Johnson et al., 2018, Kim et al., 2019] can endow conditional average treatment effect estimation with robustness to unknown covariate shifts. Indeed, the problem of confounding itself is a covariate shift problem, from the treated or control population to some target population [Johansson et al., 2022]. Multi-accurate learning is a powerful and flexible framework that, by ensuring low predictive bias over a test function class, is also robust to combinations of these covariate shifts: those induced by confounding or unknown covariate shifts in the target population. Although multi-calibrated and accurate learning originated from fairness motivations for ensuring calibration/low prediction bias over rich subgroups, in this work we show how the adversarial test functions in the formulation also confer broad robustness against covariate shift. To highlight this flexibility, we use multi-accurate calibration on an extremely small clinical trial to correct a predictor from an confounded observational study.
@@
Though there is extensive work on establishing external validity and transportability of causal effects, most of this work assumes information about a target population. Drawing inspiration from Kim et al. [2022], which studied “universal adaptability” of averaged outcome estimators with bias robust to unknown covariate shifts, we learn CATE estimates that will maintain unbiased predictions under unknown target populations.
@@
Although causal inference and machine learning has witnessed significant methodological innovation either in orthogonal/statistical learning or other machine learning adaptations [Kennedy, 2023, Nie and Wager, 2020, Chernozhukov et al., 2018, Wager and Athey, 2018, Shalit et al., 2017, Hill, 2011], to name just a few, multi-accurate learning [Kim et al., 2019] introduces a different methodological toolkit related to boosting/adversarial formulations of conditional moment conditions [Dikkala et al., 2020, Bennett and Kallus, 2023, Ghassami et al., 2022]. We conduct a thorough empirical study comparing finite- and large-sample performance of multi-accurate learning and other causal machine learning techniques more specifically tailored for causal structure. To summarize, we find that multi-accurate methods grant additional robustness against unknown covariate shifts while being competitive with more advanced causal machine learning methods in finite-samples. There is a robustness-efficiency tradeoff: the latter methods are designed to exploit in-distribution efficiency, which multi-accurate learning “off-the-shelf” does not. Nonetheless, our work connects these two previously unrelated lines of work and shows how multi-accurate learning “off-the-shelf” can address the problem of robust CATE estimation. Multi-accurate learning reduces prediction bias from model misspecification, just as is required for conditional average treatment effect estimation.
@@
In our thorough empirical study we find that our proposed multi-accurate T- and DR-learner perform well under unobserved covariate shift. Although our work does not suggest multi-accurate learning as a replacement for state-of-the-art causal machine-learning for in-distribution estimation, due to differences in variance reduction or hyperparameter selection subroutines, it does provide evidence that could inform further methodological improvements and variance reduction of multi-accurate learning for CATE estimation. In summary: Multi-accurate learning can be used “off-the-shelf” to post-process CATE estimates based on differenced outcome regressions to endow them with robustness to unknown covariate shift. Such post-processing can be appealing because it can work with only black-box access to predictors and original data. (Our earlier examples indicate many situations where only black-box access to predictors is available). Alternative approaches to robustness against unknown shifts, like distributionally robust optimization, could change the robust-optimal predictor to a risk-sensitive one rather than the true CATE, but multi-accurate learning does not.
The contributions of our work are the following.
@@
  • We propose multi-accurate post-processing of differenced-regression (T-learner) and pseudo-outcome (the doubly-robust score in the DR-learner) based CATE estimation to obtain unbiased prediction on unknown deployment populations. This approach can also flexibly adapt to a variety of covariate shifts from confounding to adversarial/unknown shifts: we illustrate by postprocessing a CATE estimator that combines large observational/small randomized data. We theoretically establish identification.
  • In Proposition 3 we show that multi-accurate post-processing of simple CATE estimates (T-learner) with a richer test function class can approximate a less-multi-accurate/less-robust but more-advanced CATE estimator, i.e. a multi-accurate DR-learner under a simpler test function class. This establishes theoretically favorable estimation properties: that our method not only provides robust bias control, but is also comparable to (weaker implementations of) CATE estimators that pursue efficient estimation. This is an important contribution since the multi-accuracy framework alone only guarantees robust bias control.
  • We show in extensive experiments with simulations and real-world observational and randomized data from the Women’s Health Initiative how our approach achieves finite-sample gains in ensuring robust bias control (and correspondingly, MSE) under unknown distribution shifts. Hence we empirically verify our theoretical contributions.
In Section 2, we introduce the formal problem setup. In Section 3 we describe the background context and most closely related work to our methodological developments (other related work is discussed in Section 6). We develop and analyze our methodology in Section 4. In Section 5, we provide extensive empirical comparisons in simulated data and a case study of the Women’s Health Initiative parallel clinical trial and observational study.
@@ 2 Problem setup
We overview the problem setup and describe directly related prior work on multi-calibration/multi-accuracy. See Section 6 for discussion of other methodological approaches.
@@
In different applications it will satisfy different assumptions, so we will later define different variants of $\mathcal{D}$. We first assume it arises from a randomized controlled trial or observational study under the assumption of weak ignorability, so that the following assumption about selection on observables holds.
@@
These are the typical so-called nuisance estimation functions used in common estimators. We use $\hat{\mu},\hat{e}$ to denote estimates.
@@
Finally, we have a mild assumption that outcomes are bounded.
Assumption 3 (Bounded outcomes).
We assume that $|Y(1)|,|Y(0)|\leq B.$
Many prior works have also assumed bounded outcomes, including but not limited to Kennedy [2020], Dudík et al. [2014]. Relaxing the assumption of bounded outcomes is a mild technicality: it is used in the convergence analysis of multi-accuracy as a sufficient but not necessary condition and as a condition for Chernoff inequalities, which could be replaced with stronger concentration inequalities under weaker assumptions. See also Devroye and Wise [1980], Gayraud [1997] for canonical estimators for estimating the support of a distribution.
@@
Later on, our work will focus on providing guarantees on conditional bias achieved by a CATE estimate $\hat{\tau}$ marginalized under a target covariate distribution $X\sim Q_{X}$ that can be different from the distribution $X\sim P$ upon which the CATE estimate was trained:
@@
We write $Q_{X}(x),P_{X},P_{X_{1}},P_{X_{0}}$ for the marginal distribution of $X$ under $Q$, the marginal distribution of $X$ under $P$, and $P_{X_{1}},P_{X_{0}}$ under $X\mid T=1,X\mid T=0$ on the observed data, respectively. We also denote ${E}_{P}[\cdot],{E}_{P_{1}}[\cdot],{E}_{P_{0}}[\cdot]$ to denote marginalization over $X$ in the training data, $X\mid T=1,$ or $X\mid T=0$, respectively. For brevity we write ${E}_{Q}[\cdot]$ to denote expectations under the unknown target distribution $Q_{X}$ on $X$ (which can be extended to accommodate shifts beyond the typical covariate shift assumption in a slight abuse of notation). For example, $\mu_{t}\in\arg\min_{g}E_{P_{X_{t}}}[(Y-g(X))^{2}],$ e.g. by default, regression in each treated arm minimizes the MSE under the covariate distribution of each treatment arm.
Figure 1: Schematic of setting 1 (external shift), and setting 2 (learning from large observational and small RCT data). We propose multi-accuracy (MC-Boost) auditing as an “off-the-shelf” procedure to improve the downstream robustness of CATE learners to unknown covariate shifts in both settings.
@@ 2.1 Robustness to unknown deployment shifts
$\textstyle\mathcal{L}_{1}$ :=$\textstyle:=\left\{\frac{dQ_{X}(x)}{dP_{X_{1}}(x)}:\mathbb{P}\left\{\frac{dQ_{X}(x)}{dP_{X_{1}}(x)}<\infty,\forall x\text{ s.t. }P(x)>0\right\}=1,\operatorname{E}[\frac{dQ_{X}(X)}{dP_{X_{1}}(X)}]=1\right\}$
$\mathcal{L}_{0}$ :=$\textstyle:=\left\{\frac{dQ_{X}(x)}{dP_{X_{0}}(x)}:\mathbb{P}\left\{\frac{dQ_{X}(x)}{dP_{X_{0}}(x)}<\infty,\forall x\text{ s.t. }P(x)>0\right\}=1,\operatorname{E}[\frac{dQ_{X}(X)}{dP_{X_{0}}(X)}]=1\right\}$
@@
$\mathcal{L}_{1}$ is the set of valid covariate shift likelihood ratios, such that $Q_{X}$ is absolutely continuous w.r.t $P_{X_{1}}$, i.e. the likelihood ratio is finite whenever $P_{X_{1}}$ has nonzero support (almost surely, with probability 1 under the measure $P_{X_{1}}$), and marginalizes to $1$. We will call this type of unknown deployment shift an “external shift”. This is analogous to the unknown shift setting studied in Kim et al. [2022], as well as other literature on unknown covariate shifts [Jeong and Namkoong, 2020, Subbaswamy et al., 2021, Hatt et al., 2021]. In contrast to an extensive literature on transportability and external validity, we focus on the case of a-priori unknown deployment shifts.
If suitably nonparametric CATE estimation indeed recovered the Bayes-optimal predictor in finite samples, there would be no issue of unknown deployment shifts. But because in finite samples it generally does not, modifying estimation to protect against unknown deployment shifts can protect against misspecification and finite-sample issues. For example, misspecified CATE estimation is vulnerable to unknown covariate shift. Furthermore, the conventional mean-squared error MSE can be nonzero for the Bayes-optimal predictor $\mu_{1}(X)={\operatorname{E}[Y(1)\mid X]}$. If the conditional bias or variance in $Y$ is heteroskedastic (i.e. varies in $x$), the prediction MSE changes as external shifts change the marginalizing covariate distribution. Under misspecification and nontrivial residual variance of the Bayes regressor, different weighting functions can indeed change the population minimizer [Shimodaira, 2000, Buja et al., 2019]. Later on we will use multi-accurate learning to post-process CATE estimates to ensure robustness against covariate shifts represented by a function class of likelihood ratios.
@@ 2.2 Unobserved confounding: observational data with RCT
We consider a different setting where unknown covariate shifts may arise: a large observational dataset and small randomized trial. The observational study may be subject to unobserved confounding. On the other hand, the sample size of the randomized data may be small, so that learning conditional causal effects solely from randomized data is unsupported. This regime is common in clinical settings, such as the parallel Women’s Health Initiative observational study and clinical trial [Machens and Schmidt-Gollwitzer, 2003]; see also [Colnet et al., 2020, Yang et al., 2020] and [Bareinboim and Pearl, 2013] for identification results for the related setting of data fusion.
@@
Assume Assumption 2. Suppose an observational dataset $\mathcal{D}_{obs}$, collected under violations of Assumption 1 (the observational data were collected under unobserved confounders), and a randomized dataset $\mathcal{D}_{rct}$, where Assumption 1 holds (the randomized data are unconfounded). For covariate shift, consider likelihood ratio functions with respect to a target population $Q$. Again we seek an estimator $\hat{\tau}(X)$ with low bias under $Q:$ $\left|\mathrm{E}_{Q}[(\hat{\tau}(X)-\tau(X))]\right|\leq\epsilon.$
@@
$|\mathcal{D}_{obs}|\gg|\mathcal{D}_{rct}|$
Assumption 4 is not necessary for identification, but it describes the relevant regime where the method is helpful: if instead $|\mathcal{D}_{rct}|\gg|\mathcal{D}_{obs}|$, unbiased CATE estimation is possible from the randomized data alone.
3 Background and related work
Conditional average treatment effect estimation.
We briefly discuss a few options for estimating $\tau$, upon which we will build multi-calibrated approaches in differing shift scenarios. The T-learner differences two regressions for the conditional means of Y for treated and untreated: